this problem is solving systems of equations. i dont know if i did the first steps correctly. the problem is:
-6y+7y=2x+2(x-2)-2x+2 (my answer was y=2x)
4(x+y)-x+y=-23 (my answer was 3x+5y=-23)
if what i did was correct, i am confused as to the next step, especially getting the first equation into an "x+y" form.

Answer by mananth(16949)

-6y+7y=2x+2(x-2)-2x+2
+y= 2x+2x-4-2x+2
y=2x-2

4(x+y)-x+y=-23
4x+4y-x+y=-23
3x+5y=-23
substitute y in second equation
3x+5y=-23
3x+5(2x-2)=-23
3x+10x-10=-23
13x=-13
x=-1
plug x
we get y = -4